Reports 1-1 of 1 Clear search Modify search
koki.okutomi - 22:51 Monday 20 May 2019 (8912) Print this report
Satellite box replacement for PRM IM OSEMs


What I did

  • I replaced the satellite boxes for the PRM IM OSEMs to modified ones.
    • The modified satellite box has additional 100 nF capacitor in the transimpedance amplification stage.
    • The old ones have been returned to AEL staffs.
    • Serial numbers of the related satellite boxes
        Old New
      IM (V1, V2, V3, H1) S1604895 S1807625
      IM (H2, H3) S1604892 S1807626
  • I measured the spectra of the IM OSEMs with both the new and old satellite boxes.
    • The new satellite boxes showed smaller noise levels at high frequencies without any malfunction.

Side works

  • I picked up the jigs for packaging the TMS-VIS from the X-end.
  • Broken traverser drivers (for PRM and PR3) were taken out of the mine.
    • These two kinds of stuffs are ready for being sent to Mitaka.

Next steps

  • Satellite box replacement for PR2 IM OSEMs
  • BO implementation for remote switching of the PR stepper motor drivers
  • Update of the information of the DGS rack configuration on the wiki
Images attached to this report
Search Help